OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. 4x^2 + ___x + 49

OpenStudy (anonymous):

this must be the square of \(2x+7\) i.e it is \[ (2x+7)^2=(2x+7)(2x+7)=4x^2+14x+14x+49=4x^2+28x+49\]

OpenStudy (anonymous):

i got 14 @satellite73